OpenStudy (anonymous):

James needs 50 tons of dirt for a large construction project. He already has 3 tons of dirt at the construction site. Each dump truck can transport 8 tons of dirt. How many trips does the dump truck need to take to reach 50 tons of dirt?

OpenStudy (vlery):

50-3=47/8=5.875 so the truck needs to go back and forth about 6 times

OpenStudy (vlery):

about 6 trips*